GEOMETRY IA What is the measure of the angle labeled (x + 12)°? (2x - 3) degrees (x + 12)
Answer #1

Here, ÁB 11 and is bansversal live. (2x-3)0 M А 4 — So, <PMA=<DNO ( A1 ternate 2 exterion 22-3 = 22 +12 2 angle) - 28-1 = 12